OpenStudy (anonymous):

Simplify (3√2 +√3)(√3 -√2)

OpenStudy (raden):

hint : (a+b)(c-d) = ac-ad+bc-bd

OpenStudy (anonymous):

Can you help me with that I dont understand you!

OpenStudy (raden):

(3√2 +√3)(√3 -√2) first, 3√2*√3 = ...

OpenStudy (anonymous):

then what I got that

OpenStudy (raden):

it should be 3√2*√3 = 3*√(2*3) = 3√6 next, 3√2 *(-√2) = ...

OpenStudy (anonymous):

which is?

OpenStudy (anonymous):

need help with that

OpenStudy (raden):

3√2 *(-√2) = -3*√4 = -3*2 = -6

OpenStudy (raden):

then, √3 * √3 = ...